tyson:

simplify -9/6 divided by 3/-2

dude:

When dividing by a fraction, just multiply by its reciprocal \(\large \frac{-9}{6}\cdot \frac{-2}{3}\)

AngeI:

Don't you have to make the bottom numbers the same? e.e

dude:

No no, thats only for addition and subtraction

AngeI:

Wut

tyson:

whats the anwser tho

dude:

Simplify \(\large \frac{-9}{6}\cdot \frac{-2}{3}\)

dude:

Only if youre doing something like \(\large \frac{9}{6}+\frac{-2}{3}\) would you need the same base

Gdeinward:

you could divide each fraction to turn it into decimal form, then just multiply the decimals.

tyson:

i just want to know what the exact anwser is

tyson:

the options are 3,1,-1,-3

dude:

Yeah... just multiply \(\large \frac{-9}{6}\cdot \frac{-2}{3}\) What is -9\(\times \)-2? What is 6\(\times \)3?

tyson:

-18 and 18

dude:

Careful, a negative times a negative is a positive Either way you do get \(\large \frac{18}{18}\) (which is just 1)

tyson:

ok thx